# On the non-Markovian quantum control dynamics

### v2

@@ V-A1 Multi-cavity case
The stability of the high-dimensional equation (73) can be similarly clarified by Proposition 9.
V-A2 An example with $M=2$
When there are two coupled cavities, the dynamics based on Markovian interaction between the quantum system and the environment has been analyzed in Ref. [82]. For the non-Markovian circumstance with $M=2$, $\mathbf{X}=\left[\langle\sigma_{-}^{(1)}\rangle,\langle\sigma_{-}^{(2)}\rangle,\langle a_{1}\rangle,\langle a_{2}\rangle\right]^{T}$, then
@@ V-A2 An example with $M=2$
$\dot{\mathbf{X}}=$ $\begin{bmatrix}-F_{1}(t)|\kappa^{(1)}|^{2}&0&-ig_{1}e^{i\delta_{1}t}&0\\ 0&-F_{2}(t)|\kappa^{(2)}|^{2}&0&-ig_{2}e^{i\delta_{2}t}\\ -ig_{1}e^{-i\delta_{1}t}&0&-\kappa&-iJ_{c}\\ 0&-ig_{2}e^{-i\delta_{2}t}&-iJ_{c}&-\kappa\end{bmatrix}\mathbf{X},$ (78)
which can be interpreted as a time-varying linear system. Similar as in Section II, we separate the vector $\mathbf{X}$ by its real and imaginary parts as $\mathbf{\mathcal{X}}=\left[\mathbf{s}^{R}_{1},\mathbf{s}^{I}_{1},\mathbf{s}^{R}_{2},\mathbf{s}^{I}_{2},\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\mathbf{a}^{R}_{2},\mathbf{a}^{I}_{2}\right]^{T}$, and denote $F_{j}(t)=F^{R}_{j}(t)+iF^{I}_{j}(t)$, which is governed by Eq. (24) after replacing $F_{R}(t)$ and $F_{I}(t)$ with $F^{R}_{j}(t)$ and $F^{I}_{j}$ respectively, then Eq. (78) can be equivalently represented with the following real-value equation as
$\dot{\mathbf{\mathcal{X}}}$ $=\begin{bmatrix}-\mathbf{\mathcal{A}}^{F}_{1}(t)|\kappa^{(1)}|^{2}&0&\mathbf{\mathcal{A}}^{G}_{1}(t)&0\\ 0&-\mathbf{\mathcal{A}}^{F}_{2}(t)|\kappa^{(2)}|^{2}&0&\mathbf{\mathcal{A}}^{G}_{2}(t)\\ \mathbf{\mathcal{A}}^{g}_{1}(t)&0&\mathbf{\mathcal{A}}_{\kappa}&\mathbf{\mathcal{A}}_{J}\\ 0&\mathbf{\mathcal{A}}^{g}_{2}(t)&\mathbf{\mathcal{A}}_{J}&\mathbf{\mathcal{A}}_{\kappa}\end{bmatrix}\mathbf{\mathcal{X}}$ (79)
$\triangleq\mathbf{\mathcal{A}}\mathbf{\mathcal{X}},$
by taking $M=2$ in Eq. (73).
Firstly, denote $\mathbf{\mathcal{X}}_{\mathbf{s}}=\left[\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\mathbf{a}^{R}_{2},\mathbf{a}^{I}_{2}\right]^{T}$, as a subspace of $\mathbf{\mathcal{X}}$. Obviously, when $g_{m}=0$ for $m=1,2$,
$\dot{\mathbf{\mathcal{X}}_{\mathbf{s}}}$ $=\begin{bmatrix}\mathbf{\mathcal{A}}_{\kappa}&\mathbf{\mathcal{A}}_{J}\\ \mathbf{\mathcal{A}}_{J}&\mathbf{\mathcal{A}}_{\kappa}\end{bmatrix}\mathbf{\mathcal{X}}_{\mathbf{s}}\triangleq\mathbf{\mathcal{A}}_{\mathbf{s}}\mathbf{\mathcal{X}}_{\mathbf{s}},$ (80)
which is linear and time-invariant. $\mathbf{\mathcal{X}}_{\mathbf{s}}$ is exponentially stable because in Eq. (78) the real part of roots determined by $(s+\kappa)^{2}+J_{c}^{2}=0$ are negative, where $s$ is for Laplace transformation. Further by Proposition 8, $\lim_{t\rightarrow\infty}\mathbf{\mathcal{X}}_{\mathbf{s}}(t)=0$. This can be further generalized to the circumstance that $M>2$.
V-B Dynamics with drives
The drive field can be applied via the feedforward or feedback channel. Take the measurement feedback as an example, as a generalization of Section IV-B, the measurement information can be collected from an arbitrary cavity of the coupled cavity array, and the measurement information from the $m$-th cavity reads
where $x_{m}=\left(a_{m}+a^{{\dagger}}_{m}\right)/\sqrt{2}$. Then we can derive the feedback dynamics according to the feedback operator similar as in Section IV.
For multiple coupled cavities, we take the feedback operator $G=\sum_{m}\left[\beta_{x}^{(m)}x+\beta_{p}^{(m)}p\right]$ with the feedback strength $g_{f}$, the dynamics of the operators can be generalized by Eq. (63) after averaging the homodyne detection noise as
$\left\langle\dot{\sigma}_{-}^{(m)}\right\rangle=-ig_{m}e^{i\delta t}\left\langle a_{m}\right\rangle-F_{m}(t)|\kappa_{1}|^{2}\left\langle\sigma_{-}^{(m)}\right\rangle,$ (82a)
$\left\langle\dot{a}_{m}\right\rangle=-i\Delta\left\langle a_{m}\right\rangle-ig_{m}e^{-i\delta t}\left\langle\sigma_{-}^{(m)}\right\rangle-\kappa\left\langle a_{m}\right\rangle$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}-ig_{f}\left(\beta_{x}^{(m)}+i\beta_{p}^{(m)}\right)\left\langle x_{m}\right\rangle-iJ_{c}\left(\left\langle a_{m-1}\right\rangle+\left\langle a_{m+1}\right\rangle\right),$ (82b)
where $\left\langle x_{m}\right\rangle=\left(\left\langle a_{m}\right\rangle+\left\langle a_{m}\right\rangle^{*}\right)/\sqrt{2}$. Obviously, the feedback drive dynamics will not directly influence the dynamics of $\left\langle\sigma_{-}^{(m)}\right\rangle$ according to Eq. (82a), but can influence the atomic dynamics via the atom-cavity couplings.
@@ V-B Dynamics with drives
When $g_{m}=0$, $\left\langle\sigma_{-}^{(m)}\right\rangle$ can stably converge to zero, while $\left\langle a_{m}\right\rangle$ can be unstable if the feedback parameter $g_{f}$ is large enough compared with the values of $F_{j}(t)$ and $\kappa$.
Denote $X_{u}=\left[\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\cdots,\mathbf{a}^{R}_{M},\mathbf{a}^{I}_{M}\right]$, and $X_{s}=\left[\mathbf{s}^{R}_{1},\mathbf{s}^{I}_{1},\cdots,\mathbf{s}^{R}_{M},\mathbf{s}^{I}_{M}\right]$, then Eq. (82) can be equivalently written as
@@
$\begin{bmatrix}\dot{X}_{u}\\ \dot{X}_{s}\end{bmatrix}$ $=\left(\begin{bmatrix}\mathbf{\mathcal{A}}^{u}&0\\ 0&\mathbf{\mathcal{A}}^{s}\end{bmatrix}+\begin{bmatrix}P_{1}(t)&P_{2}(t)\\ P_{3}(t)&P_{4}(t)\end{bmatrix}\right)\begin{bmatrix}X_{u}\\ X_{s}\end{bmatrix},$ (83)
where
$\mathbf{\mathcal{A}}^{u}=$ $\begin{bmatrix}\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa&\Delta&\cdots&0&0\\ \sqrt{2}g_{f}\beta_{x}^{(1)}-\Delta&-\kappa&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\sqrt{2}g_{f}\beta_{p}^{(M)}-\kappa&\Delta\\ 0&0&\cdots&\sqrt{2}g_{f}\beta_{x}^{(M)}-\Delta&-\kappa\end{bmatrix}$ (84)
$\mathbf{\mathcal{A}}^{s}$ $=-\begin{bmatrix}\bar{R}_{f}^{(1)}&-\bar{I}_{f}^{(1)}&\cdots&0&0\\ \bar{I}_{f}^{(1)}&\bar{R}_{f}^{(1)}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\bar{R}_{f}^{(M)}&-\bar{I}_{f}^{(M)}\\ 0&0&\cdots&\bar{I}_{f}^{(M)}&\bar{R}_{f}^{(M)}\end{bmatrix},$ (85)
$\mathbf{\mathcal{A}}_{J}$ is firstly given by Eq. (77), $\bar{R}_{f}^{(m)}=\lim_{t\rightarrow\infty}F^{R}_{m}(t)$, $\bar{I}_{f}^{(m)}=\lim_{t\rightarrow\infty}F^{I}_{m}(t)$, $m=1,2,\cdots,M$, $P_{1}(t)=0$, $P_{2}(t)=\mathbf{R}(t)$, $P_{3}(t)=\mathbf{G}(t)$, $P_{4}(t)=\mathbf{F}(t)-\mathbf{\mathcal{A}}^{s}$, $\mathbf{R}(t)$, $\mathbf{G}(t)$ and $\mathbf{F}(t)$ are given by Eqs. (76), (75), and (74), respectively.
For a special case that $g_{m}=0$, then $P_{2}(t)=P_{3}(t)=0$, the dynamics of $X_{u}$ and $X_{s}$ are decoupled in Eq. (83). Obviously, $X_{s}$ will converge to zero if only the interaction between the atom and the environment becomes Markovian. However, $X_{u}$ can be unstable because of the feedback.
When $g_{m}\neq 0$, $X_{u}$ and $X_{s}$ are always coupled, the stability of Eq. (83) can be determined by the following proposition.
Proposition 11.
When the interaction between the atom and the environment becomes asymptotically Markovian, $X_{u}$ approaches zero when $t\rightarrow\infty$ provided that
$\int_{t_{0}}^{t_{0}+T}\mu\begin{Bmatrix}\begin{bmatrix}\mathbf{\mathcal{A}}^{u}&P_{2}(t)\\ P_{3}(t)&\mathbf{\mathcal{A}}^{s}\end{bmatrix}\end{Bmatrix}\mathrm{d}\tau<0.$ (86)

### v3

@@ V-B Dynamics with drives
When $g_{m}=0$, $\left\langle\sigma_{-}^{(m)}\right\rangle$ can converge to zero, while $\left\langle a_{m}\right\rangle$ can be unstable if the feedback parameter $g_{f}$ is large enough compared with $\left|\mathcal{F}_{m}(t)\right|$ and $\kappa$.
Denote $X_{u}=\left[\mathbf{a}^{R}_{1},\mathbf{a}^{I}_{1},\cdots,\mathbf{a}^{R}_{M},\mathbf{a}^{I}_{M}\right]$, and $X_{s}=\left[\mathbf{s}^{R}_{1},\mathbf{s}^{I}_{1},\cdots,\mathbf{s}^{R}_{M},\mathbf{s}^{I}_{M}\right]$. Then Eq. (70) can be equivalently written as
$\begin{bmatrix}\dot{X}_{u}\\ \dot{X}_{s}\end{bmatrix}$ $=\left(\begin{bmatrix}\mathbf{\mathcal{A}}_{u}&0\\ 0&\mathbf{\mathcal{A}}_{s}\end{bmatrix}+\begin{bmatrix}\mathcal{Q}_{1}(t)&\mathcal{Q}_{2}(t)\\ \mathcal{Q}_{3}(t)&\mathcal{Q}_{4}(t)\end{bmatrix}\right)\begin{bmatrix}X_{u}\\ X_{s}\end{bmatrix},$ (71)
where
@@
$\mathbf{\mathcal{A}}_{u}=$ $\begin{bmatrix}\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa&\Delta&\cdots&0\\ \sqrt{2}g_{f}\beta_{x}^{(1)}-\Delta&-\kappa&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\Delta\\ 0&0&\cdots&-\kappa\end{bmatrix}$ (72)
$\mathbf{\mathcal{A}}_{s}$ $=-\begin{bmatrix}\bar{R}_{f}^{(1)}&-\bar{I}_{f}^{(1)}&\cdots&0&0\\ \bar{I}_{f}^{(1)}&\bar{R}_{f}^{(1)}&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\bar{R}_{f}^{(M)}&-\bar{I}_{f}^{(M)}\\ 0&0&\cdots&\bar{I}_{f}^{(M)}&\bar{R}_{f}^{(M)}\end{bmatrix},$ (73)
$\mathbf{\mathcal{A}}_{J}$ is given in Eq. (67), $\bar{R}_{f}^{(m)}=\lim_{t\rightarrow\infty}\mathcal{F}^{R}_{m}(t)$, $\bar{I}_{f}^{(m)}=\lim_{t\rightarrow\infty}\mathcal{F}^{I}_{m}(t)$, $m=1,2,\cdots,M$, $\mathcal{Q}_{1}(t)=0$, $\mathcal{Q}_{2}(t)=\mathbf{R}(t)$, $\mathcal{Q}_{3}(t)=\mathbf{G}(t)$, $\mathcal{Q}_{4}(t)=\mathbf{F}(t)-\mathbf{\mathcal{A}}_{s}$, $\mathbf{R}(t)$, $\mathbf{G}(t)$ and $\mathbf{F}(t)$ are given by Eqs. (66), (65), and (64), respectively.
For a special case that $g_{m}=0$, $P_{2}(t)=P_{3}(t)=0$, the dynamics of $X_{u}$ and $X_{s}$ are decoupled in Eq. (71). Obviously, $X_{s}$ will converge to zero if only the interaction between the atom and the environment becomes Markovian. However, $X_{u}$ can be unstable because of the feedback.
When $g_{m}\neq 0$, $X_{u}$ and $X_{s}$ are always coupled, the stability of Eq. (71) can be determined by the following theorem based on the function $\mu(\cdot)$ given by Definition 4 in Appendix C.
Theorem 3.
When the interaction between the atom and the environment becomes asymptotically Markovian in Eq. (71), $X_{u}$ approaches zero when $t\rightarrow\infty$ provided that
$\int_{t_{0}}^{t_{0}+T}\mu\begin{Bmatrix}\begin{bmatrix}\mathbf{\mathcal{A}}_{u}&\mathcal{Q}_{2}(\tau)\\ \mathcal{Q}_{3}(\tau)&\mathbf{\mathcal{A}}_{s}\end{bmatrix}\end{Bmatrix}\mathrm{d}\tau<0.$ (74)
Proof.
When the interaction between the atom and the environment becomes asymptotically Markovian, $\lim_{t\rightarrow\infty}\mathcal{Q}_{4}(t)=0$ in Eq. (71) with $\mathcal{Q}_{1}(t)=0$, and the matrix $\begin{bmatrix}\mathbf{\mathcal{A}}_{u}&\mathcal{Q}_{2}(t)\\ \mathcal{Q}_{3}(t)&\mathbf{\mathcal{A}}_{s}\end{bmatrix}$ is periodic. Then the result holds according to the proof of Corollary 2. ∎
@@
As in Ref. [59], a choice of the norm in Theorem 3 can be
@@
$\frac{1}{2}\lambda_{\max}\left(\begin{bmatrix}\mathbf{\mathcal{A}}_{u}+\mathbf{\mathcal{A}}_{u}^{\rm T}&\mathcal{Q}_{2}(t)+\mathcal{Q}_{3}^{\rm T}(t)\\ \mathcal{Q}_{3}(t)+\mathcal{Q}_{2}^{\rm T}(t)&\mathbf{\mathcal{A}}_{s}+\mathbf{\mathcal{A}}_{s}^{\rm T}\end{bmatrix}\right)$ (75)
where $\bar{\mathbf{a}}_{m}=\begin{bmatrix}2\sqrt{2}g_{f}\beta_{p}^{(m)}-2\kappa_{m}&\sqrt{2}g_{f}\beta_{x}^{(m)}\\ \sqrt{2}g_{f}\beta_{x}^{(m)}&-2\kappa_{m}\end{bmatrix}$, $\bar{\mathbf{b}}_{m}=\begin{bmatrix}-2\bar{R}_{f}^{(m)}&0\\ 0&-2\bar{R}_{f}^{(m)}\end{bmatrix}$. When $g_{f}$ and $g_{x}$ are small, the inequality in (74) can be easier to be satisfied, then both $X_{u}$ and $X_{s}$ can converge to zero. An extreme case is that $\beta_{x}^{(m)}=0$ and $\kappa_{m}>\sqrt{2}g_{f}\beta_{p}^{(m)}$, then the matrix in Eq. (LABEL:con:MPMtr) is a diagonal matrix with all the diagonal elements being negative.
As compared in Fig. 6, there are two coupled cavities, $J_{c}=0.1$GHz, $\delta=0.2$GHz, $\kappa=1$GHz, $g_{1}=0.2$GHz, $g_{2}=0.4$GHz, $\Delta=5$GHz, the two atoms in the two cavities are coupled identically with the non-Markovian environment with $\omega_{a}=43.98$GHz, and $\Omega=38.98$GHz, $\chi=1$, $\kappa^{(1)}=\kappa^{(2)}=0.04$GHz $\gamma=2$GHz for the non-Markovian environment [53]. In the measurement feedback design, $\beta_{x}^{(1)}=\beta_{x}^{(2)}=0$GHz, $\beta_{p}^{(1)}=\beta_{p}^{(2)}=0.2$GHz. On one hand, when $g_{f}=1$, $\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa=-0.7172{\rm GHz}<0{\rm GHz}$. On the other hand, when $g_{f}=7$, $\sqrt{2}g_{f}\beta_{p}^{(1)}-\kappa=0.9799{\rm GHz}>0{\rm GHz}$. The two group of simulations agree with Remark 7, showing that the stable and unstable subspaces can be modulated by tuning the feedback parameters.
The above analysis on multiple Jaynes-Cummings model realized by cascade cavity networks can be a potential approach for realizing non-Markovian interactions, according to the introduction on the frequency spectrum properties in [71]. Besides, the coupled cavity array can be further used to realize waveguide quantum electrodynamics [72], constructing various complex non-Markovian quantum networks [10, 11, 13]. The above generalization can introduce more challenges on quantum control due to the fact that the non-Markovian environment can destruct the purity of quantum states [73], which can further influence the control properties such as the creation of entangled quantum states [20] and photons [13].
VI Conclusion
In this paper, we study quantum control dynamics within a cavity-QED system, focusing on the non-Markovian interactions between the quantum system and its environment. The evolution of parameters representing the atom’s non-Markovian decay into the environment is described by nonlinear dynamics. The transition to Markovian interactions between the quantum system and the environment is explained by the stability of these nonlinear processes. Consequently, the dynamics of the multi-level system in a non-Markovian environment can be modeled by linear time-varying equations. Manipulation of atomic states and photons in the cavity is then achieved using open-loop and closed-loop control methods that utilize quantum measurement feedback. This approach can be extended to scenarios involving multiple coupled cavities described by high-dimensional linear time-varying equations, where feedback control can further influence the dynamics between stable and unstable subspaces.
Appendix A Derivation of the non-Markovian parameter dynamics
We take the atom and cavity as a whole represented with the state $|\psi(t)\rangle$, then the combined system interacts with the bath via the multi-level atom. The interaction between the atom and environment can be represented with the Hamiltonian
with the detailed meaning introduced after Eq. (2) in the main text. Here we take a representative simplified case that the quantum system is coupled to the environment via operators $L_{n}$. The evolution of quantum states is influenced by its stochastic interaction with the environment as [27, 26, 41, 42]
$\frac{\delta|\psi(t)\rangle}{\delta z_{s}}=\sum_{n=1}^{N-1}f_{n}(t,s)L_{n}|\psi(t)\rangle,$ (76)
where $z_{s}$ is a complex-valued Wiener process at time $s$, $f_{n}(t,s)$ is a two-time function to be determined, related to the two-time points $t$ and $s$ [27].
Combined with Eq. (3), the differential of Eq. (76) upon the time $t$ reads
$\frac{\delta|\dot{\psi}\rangle}{\delta z_{s}}=\sum_{n=1}^{N-1}\left[\frac{\partial f_{n}(t,s)}{\partial t}L_{n}|\psi\rangle+f_{n}(t,s)L_{n}|\dot{\psi}\rangle\right],$ (77)
where $|\dot{\psi}\rangle$ represents the derivation according to time $t$, and
$\frac{\partial f_{n}(t,s)}{\partial t}L_{n}|\psi\rangle=\frac{\delta|\dot{\psi}\rangle}{\delta z_{s}}-f_{n}(t,s)L_{n}|\dot{\psi}\rangle$ (78)
$=-iH\frac{\delta|\psi\rangle}{\delta z_{s}}+if_{n}(t,s)L_{n}H|\psi\rangle+L_{n}\frac{\delta|\psi\rangle}{\delta z_{s}}z_{t}$
$~{}~{}~{}-\frac{\delta}{\delta z_{s}}L_{n}^{{\dagger}}\int_{0}^{t}\alpha(t,s)\frac{\delta|\psi\rangle}{\delta z_{s}}\mathrm{d}s-f_{n}(t,s)L_{n}L_{n}|\psi\rangle z_{t}$
$~{}~{}~{}+f_{n}(t,s)L_{n}L_{n}^{{\dagger}}\int_{0}^{t}\alpha(t,s)\frac{\delta|\psi\rangle}{\delta z_{s}}\mathrm{d}s$
$=i\left[L_{n},H\right]f_{n}(t,s)|\psi\rangle$
$~{}~{}~{}+F_{n}(t)\left(L_{n}L_{n}^{{\dagger}}-L_{n}^{{\dagger}}L_{n}\right)f_{n}(t,s)L_{n}|\psi\rangle,$
with the time-dependent function $F_{n}(t)$ defined by Eq. (7) and $\alpha(t,s)$ given by Eq. (5) in the main text.
Then combined with Eq. (3), for a specific $L_{n}$, Eq. (78) can be rewritten as
$\frac{\partial f_{n}(t,s)}{\partial t}L_{n}|\psi\rangle$ (79)
$=$ $i\omega_{n}f_{n}(t,s)L_{n}|\psi\rangle+F_{n}(t)\kappa_{n}f_{n}(t,s)L_{n}|\psi\rangle,$
further combined with Eq. (7), we have
$\dot{F}_{n}(t)=\frac{\gamma\chi_{n}}{2}-(\gamma+i\Omega)F_{n}(t)$ (80)
$~{}~{}~{}~{}+\int_{0}^{t}\alpha_{n}(t,s)\left[i\omega_{n}f_{n}(t,s)+F_{n}(t)\kappa_{n}f_{n}(t,s)\right]\mathrm{d}s$
$=\kappa_{n}F_{n}^{2}(t)-\left(\gamma+i\Omega-i\omega_{n}\right)F_{n}(t)+\frac{\gamma\chi_{n}}{2},$
where $f_{n}(t,t)=\chi_{n}$ is a constant, $F(0)=0$, and the analytic solutions of generalized $F_{n}(t)$ is given in Ref. [27].
Appendix B Derivation of the non-Markovian master equation
The non-Markovian master equation has been introduced in Refs. [27, 29]. In this Appendix, we briefly introduce the derivation of the non-Markovian control equation (6) in the main text, which is the control equation we focus on in this paper.
The evolution of $f_{n}(t,s)$ has been clarified in Appendix A. Combining Eq. (4) with Eq. (76) after the normalization, we have the following equation for the state vector,
$\frac{\mathrm{d}}{\mathrm{d}t}|\psi(t)\rangle=-iH|\psi(t)\rangle+\sum_{n=1}^{N-1}\left(L_{n}-\left\langle L_{n}\right\rangle\right)|\psi(t)\rangle z_{t}$ (81)
$-\sum_{n=1}^{N-1}\left(L_{n}^{{\dagger}}-\left\langle L_{n}^{{\dagger}}\right\rangle\right)\int_{0}^{t}\alpha(t,s)f_{n}(t,s)L_{n}|\psi(t)\rangle\mathrm{d}s.$
Considering that $\rho(t)=E_{z}(|\psi(t)\rangle\langle\psi(t)|)$, and denote $P_{t}\left(z_{t}^{*}\right)=|\psi(t)\rangle\langle\psi(t)|$, then $E_{z}\left(P_{t}z_{t}\right)=\sum_{n=1}^{N-1}\int_{0}^{t}E_{z}\left[z_{t}z_{s}^{*}\right]f_{n}^{*}(t,s)\rho(t)L_{n}^{{\dagger}}\mathrm{d}s$ and $E_{z}\left(P_{t}z_{t}^{*}\right)=\sum_{n=1}^{N-1}\int_{0}^{t}E_{z}\left[z_{t}^{*}z_{s}\right]f_{n}(t,s)L_{n}\rho(t)\mathrm{d}s$ [42]. Thus the dynamics of the density matrix reads
$\dot{\rho}=-i[H,\rho]+\sum_{n=1}^{N-1}\left\{\int_{0}^{t}\mathrm{d}sE_{z}\left[z_{t}z_{s}^{*}\right]f_{n}^{*}(t,s)+\right.$ (82)
$\left.\int_{0}^{t}\mathrm{d}sE_{z}\left[z_{t}^{*}z_{s}\right]f_{n}(t,s)L_{n}\rho L_{n}^{{\dagger}}-\int_{0}^{t}\alpha(t,s)f_{n}(t,s)\mathrm{d}s\right.$
$\left.L_{n}^{{\dagger}}L_{n}\rho-\rho L_{n}^{{\dagger}}L_{n}\int_{0}^{t}\alpha^{*}(t,s)f_{n}^{*}(t,s)\mathrm{d}s\right\},$
Appendix C Some results on nonlinear and linear dynamics
In this appendix, we recall some concepts and conclusions on nonlinear and linear time-varying systems for the clarifications in Sec. II and Sec. III.
Definition 1 ([74, 75]).
The system in Eq. (28) is said to be bounded-input bounded-output (BIBO) stable if for every $a_{u}\geq 0$ and $\alpha_{X}\geq 0$ there is a finite number $\beta_{X}=\beta_{X}\left(a_{u},\alpha_{X}\right)$ such that for every initial condition $X_{F}\left(t_{0}\right)$ with $X_{F}\left(t_{0}\right)\leq\alpha_{X}$ and every sequence $u$ with $\|u\|\leq a_{u}$,
for $t^{\prime}>t_{0}$.
Definition 2 ([51]).
For a state vector $X_{F}(t)$ evolves in the time domain as $\dot{X}_{F}(t)=\mathcal{F}\left(X_{F}\right)$, a set $S$ is said to be forward invariant if $X_{F}\left(t_{0}\right)\in S$, $X_{F}\left(t\right)\in S$ for any $t>t_{0}$.
Proposition 5 ([74]).
Suppose for each $u\geq 0$, there exists a positive Lyapunov function $V_{u}$ for the system with the input $u$ and
$\dot{V}_{u}(t,X_{f})\leq 0.$
When the input satisfies $\|u\|\leq a_{u}$, then the system is BIBO stable.
Proposition 6 ([76]).
For the nonlinear system
Definition 3 ([59]).
$\|\Phi\left(t,\tau\right)\|\leq Ke^{-\tilde{\alpha}\left(t-\tau\right)},$
for $t_{0}\leq\tau<t<\infty$.
Definition 4 ([59]).
$\mu[A(t)]=\lim_{h\rightarrow 0^{+}}\frac{\|I_{n}+hA(t)\|-1}{h},$
and
$\Pi^{+}(t)\triangleq\int_{t_{0}}^{t}\mu[A(\tau)]\mathrm{d}\tau.$ (83)
Appendix D Proof of Corollary 2